solving equations using several operations. don't understand fractions when it comes to equations an inequalities. need help. a/3+1=7.
Answer by checkley71(8403)

a/3+1=7 or a/3=7-1 or a/3=6 or a=18
proof 18/3+1=7 or 6+1=7 or 7=7